calculate the percentage of p-charater in the orbitals of P4  forming p-p bond

Grade:11

6 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Hybridization of P4 is sp3.
which means one s-orbital and 3 p-orbitals.
% p = (p-orbitals / total orbitals)x100
%p = (3/4)x100 = 75 %

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

P is SP3 hybridised in P4  so P character is 75%

NEERAJ DUBEY
13 years ago
in p4 phosphorus is p3 hybridised s orbotal do not participate on bonding lone pair electrons are prsent in pure s orbital thus 100% p charecter
in p4 molecule s orbital do not participate in bonding so molecule is p3 hybridised so p char. 100%
Satish
6 years ago
There is no sp3 hybridisation in p4Bcoz there is participate pure p orbital in hybridisation Sooo there is 100% pure p orbital
Deepali
5 years ago
In P4 molecule there are all single bonds which means that  P4 has sp3 hybridization. Hence % p character = 3/4*100 = 75% which is the required answer. 
Raghav
4 years ago
Their is no hybridization in P4 due to Drago's rule ( en of p is 2.12 and p is of 3rd period) and it's lone pair is in inactive s orbital just like PH3